1. What is the total charge of the neptunium nucleus? (The neutral neptunium atom has 93 electrons.)

a. What is the magnitude of its electric field at a distance of 6.76×10-10 m from the nucleus?

b. What is the magnitude of the force on an electron at that distance?

c. What would the magnitude of the force be if the distance of the electron from the nucleus were doubled?

Use all charges in Coulombs
Answer #1

a)

total charge = 93* 1.6* 10^-19 = 1.488* 10^-17 C

b)

E = k q/r^2 = 9* 10^9* 1.488* 10^-17 / (6.76* 10^-10)^2

E = 2.93* 10^11 N/C

c)

F = eE = 4.689* 10^-8 N

d)

as we know force is inversely proportional to square of the distance between tthe charges

so, if distance doubled, force decreased to (1/4) of its value

F' = F/4 = 1.172* 10^-8 N

=======

Comment before rate in case any doubt, will reply for sure.. goodluck